# IDEAL: Influence-Driven Selective Annotations Empower In-Context Learners in Large Language Models

### v1

@@ 1 Introduction
In-context learning (ICL) entails presenting a small set of examples with demonstrations as prompts (called in-context examples) to large language models (LLMs), before making predictions on test inputs [1, 2, 3]. This emerging few-shot learning paradigm is an appealing alternative to supervised fine-tuning as it can avoid heavy parameter updates of language models while improving accuracy [4, 5].
Recent studies indicate that obtaining prompts from a vast collection of annotated examples is crucial to achieving strong performance [6]. Notably, these studies have illuminated the substantial performance improvements when retrieving analogous examples (under specific embedding criteria) as in-context examples tailored for each individual test input. Since different test scenarios need distinct in-context examples, and each of them is equipped with its pertinent annotations, the necessity of a large volume of annotated examples is emphasized [7]. However, obtaining large-scale annotated examples for ICL requires substantial manpower and financial resources. This is because humans not only need to annotate the true label for each example but also need to provide the example demonstration in the annotation process [8].
@@
Figure 1: Visualization of the information diffusion process [9] of two subsets with equal sizes. Experiments are conducted using the SST-5 training set [10]. To avoid the denseness, we randomly sample 100 examples in total. In this visualization, black nodes present the initial subset without information diffusion. White nodes correspond to the examples that are not influenced by diffusion. For other nodes, darker nodes represent earlier influenced examples. We can observe that the subset with high influence (b) can achieve better performance by influencing a larger group of examples in the unlabeled data pool compared to the subset with low influence (a).
@@
To reduce the annotation cost, the previous effort Vote-$k$ [7] made attempts by proposing to select a diverse and representative subset from a large-scale unlabeled data pool to annotate. Particularly, Vote-$k$ initially selects a small portion of data for diversity and annotates them manually. Then, these annotated data act as prompts for predictions on all other unlabeled data, and choose the remaining ones that need to be annotated, based on diverse confidence scores. However, despite its pretty performance in empirical evaluations, Vote-$k$ is still unsatisfactory in practice. We detail the issues from three aspects. (1) The data selection procedure of Vote-$k$ is not end-to-end. This results in inconvenience, increased processing complexity, and added inference costs due to the predictions on unlabeled data. (2) Diversity and representativeness need to be balanced carefully [7]. Highlighting diversity in data selection is crucial for comprehensive coverage, but may sacrifice representativeness by overlooking exemplary data. Besides, the excessive emphasis on diversity of Vote-$k$ causes the selection of outliers (see evidence in Appendix C.2). (3) Vote-$k$ devoids theoretical guarantees, making it challenging to assess the algorithm’s reliability in realistic tasks and constraining its practical utility.
@@
In this paper, to minimize annotation costs for ICL and address the issues of existing work, an innovative data selection method is introduced, where we utilize influence-driven selective annotations to empower in-context learners (IDEAL). In essence, IDEAL aims to identify a subset of data that acts as a proxy and closely approximates the vast unlabeled dataset. Once annotated, these selected data can be considered a viable substitute for the large annotated examples in subsequent ICL tasks. In further detail, our method works in an unsupervised and end-to-end manner. We first construct a directed graph, where its vertices represent unlabeled data and its edges bridge different data based on their similarities. Inspired by influence maximization that aims to select a vertex set at key positions in social graphs [11], we then propose to quantify the influence of each candidate unlabeled subset in our constructed graph, through a classic independent-cascade diffusion model illustrated in Figure 1. To find the subset with high influence, a simple greedy algorithm for unlabeled data selection is introduced. The algorithm does not need a delicate trade-off between diversity and representativeness. Instead, it iteratively selects a vertex if it provides a maximum marginal gain to the influence metric, until the selection is completed based on the annotation budget.
@@
Theoretically, under the influence-driven selective paradigm, we provide the lower bound for the subset influence selected by our method, demonstrating it is at least as large as a certain proportion of the influence of the optimal solution. Empirically, we conduct comprehensive experiments over 9 datasets across diverse tasks (covering classification, commonsense reasoning, dialogue, and text/code generation). Various LLMs and prompt retrieval technologies are included in evaluations. Experimental results demonstrate that our IDEAL can achieve better performance than Vote-$k$, with only 13% time consumption during subset selection. This creates a strong baseline of selective annotations for follow-up research.
@@ 2.2 Influence-driven selective annotations
Overview. For selective annotations in ICL, we need to identify a subset that approximates vast unlabeled data. Therefore, quantifying the coverage of each candidate subset is critical. To achieve this, we construct a directed graph using the embeddings of unlabeled data and portray their relationships using the edges in the graph. We then quantify the influence of each candidate subset in the constructed graph. An information diffusion model is used for this purpose. After the quantification, we can search the subset with maximum influence, which most closely approximates the unlabeled data. Below we detail the above procedure step by step.
Constructing the directed graph. We first compute a vector embedding for each unlabeled instance using Sentence-BERT [12] [^1] https://huggingface.co/sentence-transformers/all-mpnet-base-v2.. The obtained embeddings are employed to build a directed graph $\mathcal{G}=(\mathbf{V},\mathbf{E},\mathbf{P})$, where the vertices $\mathbf{V}=\{\mathbf{v}_{i}\}_{i=1}^{n}$ represent the embeddings of the unlabeled instances, $\mathbf{E}$ denotes the set of edges in the graph, and $\mathbf{P}$ denotes the set of weights assigned to edges. In more detail, for each vertex $\mathbf{v}\in\mathbf{V}$, we connect it to its $k$ nearest successors [^2] In graph theory [13], a vertex $\mathbf{u}$ is the successor of a vertex $\mathbf{v}$ if it is at the end of an outgoing directed edge ($\mathbf{v}$,$\mathbf{u}$). in terms of the cosine similarity between the embeddings and then get $\mathbf{E}$. For the edge $(\mathbf{v},\mathbf{u})\in\mathbf{E}$ that connects $\mathbf{v}$ and its successor $\mathbf{u}$, we assign the weight $p(\mathbf{v},\mathbf{u})=\cos(\mathbf{v},\mathbf{u})/{\sum_{\mathbf{z}\in\mathcal{N}(\mathbf{v},k)}\cos(\mathbf{v},\mathbf{z})}$ with $p\in\mathbf{P}$, where $\mathcal{N}(\mathbf{v},k)$ represents the set including $k$ nearest successors of $\mathbf{v}$, and $\cos(\cdot,\cdot)$ is a function that calculates the cosine similarity of two embeddings. The constructed graph depicts the relationships between unlabeled examples in terms of the embedding similarity.
@@
At the beginning, the subset $\mathcal{S}$ is activated. Then at each step, each vertex $\mathbf{v}$ activates its successors that remained inactive in the last step with a probability defined by $p(\mathbf{v},\mathbf{u})$. The activation can be conceptualized as a coin toss where the outcome is determined by the head probability $p(\mathbf{v},\mathbf{u})$. If the result is the head, the vertex $\mathbf{v}$ becomes activated; otherwise, it remains inactive. Starting from $\mathcal{S}$, the diffusion terminates when no further vertex can be activated in the graph. Lastly, we quantify the influence of the set with the number of activated vertices, where a larger number corresponds to greater influence. To help understand the procedure of Algorithm 1, we provide an illustration as shown in Figure 2. For convenience, we express Algorithm 1 as an influence function $f_{\mathcal{G}}(\mathcal{S})$ for the graph $\mathcal{G}$ that takes example set $\mathcal{S}$ as inputs, and returns the number of activated vertices $L$.
@@
Searching the subset with maximum influence. We exploit a simple yet effective greedy algorithm [14] to search the subset with maximum influence, which is illustrated in Algorithm 2. Specifically, the algorithm is initialized with an empty set, and iteratively involves an instance if it can provide the maximum marginal gain to the influence function. The search algorithm terminates when the selected subset meets the annotation budget. Finally, we achieve the set $\mathcal{S}_{\rm{u}}$ that includes $m$ examples to annotate, using the correspondence between embeddings and instances.
@@ 2.3 Prompt retrieval
After the above influence-driven selective annotations, the subset $\mathcal{S}_{\rm{u}}$ is achieved. By making manual annotations on $\mathcal{S}_{\rm{u}}$, a set of annotated examples is obtained. We can then retrieve examples from the annotated set as in-context examples for each test input. Following previous studies [4, 7], we will calculate embeddings for all annotated examples using Sentence-BERT [12] and identify the most similar instances to each test input based on the cosine similarity. Notice that, the proposed method is agnostic to prompt retrieval methods. As demonstrated in §4.3.3, our method can be combined with any other prompt retrieval technologies. Better prompt retrieval technologies can further boost final performance.
@@ 3 Theoretical Analysis
Remark 1. Intuitively speaking, given any graph $\mathcal{G}$, we say the influence function $f_{\mathcal{G}}$ satisfies the submodular condition if adding one data point to a smaller subset provides more influence than adding the same data point to a larger subset. In other words, it reflects the principle of diminishing returns: the marginal gain of including a data point in a set decreases as the size of the set increases. This condition can hold within the influence function [15]. Considering an extreme case, when subset $\mathcal{S}=\mathbf{V}$, the influence improvement of adding any data point to $\mathcal{S}$ will be zero.
@@ 4.1 Experimental setups
Datasets and tasks. Following previous work [7], we employ 9 datasets for the evaluations, which can be categorized into 4 different tasks, including classification, multi-choice, dialogue, and generation. The details of the datasets are provided in Appendix D.1. For each dataset, the original “train/dev/test” split from the Transformer library [16] is utilized. We use test data for evaluation if they are available publicly (SST-5 [10], DBpedia [17], MWoZ [18], and Xsum [19]). Otherwise, we follow the same setting in [7] and use the development set. We use accuracy as metric for all classifications and multiple choices tasks, joint accuracy [18] for MWoZ, test suite accuracy [20] for GeoQuery [21], and ROUGE-L [22] for Xsum.
Models. If not otherwise specified, we run all experiments on the GPT-J 6B model [23] except the GeoQuery and MWoZ datasets where we use Text-devinci-002 [24]. We also provide experiments on other models including GPT-Neo 2.7B [25] and more advanced models GPT-3.5-Turbo [26] in §4.3.4. Our implementation is detailed in Appendix D.2.
Baselines. In the main experiments, we perform a comprehensive evaluation of our method that is compared with previous state-of-the-art selective annotation baselines, i.e., Vote-$k$ [7] and random selection (abbreviated as “Random” below). Note that, in §4.3.2, we also compare our method with alternative methods that can select a coreset from large-scale unlabeled data on typical datasets. For the baseline Vote-$k$, we conduct experiments by running its official code [^3] https://github.com/HKUNLP/icl-selective-annotation..
@@ 4.2 Main results
Measurement on performance. We first perform the evaluations for Random, Vote-$k$, and our method. The annotation budget is set to 18 and 100 respectively. Note that we include $18$ as the annotation budget considering all annotated examples can be fit to the prompt of the large language models within context limits. Therefore, the prompt retrieve stage can be ignored and the evaluation results can naturally represent the quality of the selected examples. We provide experimental results in Table 1. As can be seen, our method achieves better performance than baselines in most of the evaluation scenarios (17 out of 18). Interestingly, we find that random selection outperforms Vote-$k$ in 3 out of 18 cases. We conjecture that, under some ideal circumstances, the selected subset by random selection can approximate the distribution of full data. If test data follows the same distribution, good performance can be achieved. Note that we also illustrate selected examples and label distributions in selective annotations in Appendix C.1 and Appendix C.4 respectively.
Measurement on time cost. Previous work Vote-$k$ [7] encompasses generating prediction for most unlabeled data with a set of selected examples as prompts and performs data selection according to the confidence scores of the prediction. However, this process results in large unnecessary costs at inference time. Meanwhile, LLMs are often used as a service and an extra charge will appear with the usage of the token in both the input and output. In Figure 3, we compare the time cost of subset selection in our method against Vote-$k$ on all tasks with the same hardware. The annotation budget is set to 18. We can observe that our method saves a tremendous amount of cost compared to Vote-$k$. Specifically, under the same hardware conditions, IDEAL achieves a 7.8$\times$ lead on average over Vote-$k$. The speed improvement benefits from the fact that the proposed method does not need to perform example selection by generating predictions on a large number of unlabeled examples and is completely unsupervised.

### v2

@@ 1 Introduction
In-context learning (ICL) entails presenting a small set of examples with demonstrations as prompts (called in-context examples) to large language models (LLMs), before making predictions on test inputs (Wei et al., 2022a; Min et al., 2022; Akyürek et al., 2023). This emerging few-shot learning paradigm is an appealing alternative to supervised fine-tuning as it can avoid heavy parameter updates of language models while improving accuracy (Liu et al., 2021; Yoo et al., 2022).
Recent studies indicate that obtaining prompts from a vast collection of annotated examples is crucial to achieving strong performance (Rubin et al., 2022). Notably, these studies have illuminated the substantial performance improvements when retrieving analogous examples (under specific embedding criteria) as in-context examples tailored for each individual test input. Since different test scenarios need distinct in-context examples, and each of them is equipped with its pertinent annotations, the necessity of a large volume of annotated examples is emphasized (Su et al., 2023). However, obtaining large-scale annotated examples for ICL requires substantial manpower and financial resources (Baldridge & Osborne, 2004; Engelson & Dagan, 1996; Snow et al., 2008). This is because humans not only need to annotate the true label for each example but also need to provide the example demonstration in the annotation process (Wei et al., 2022b).
@@
Figure 1: Visualization of the information diffusion process (Goldenberg et al., 2001) of two subsets with equal sizes. Experiments are conducted using the SST-5 training set (Socher et al., 2013). To avoid the denseness, we randomly sample 100 examples in total. In this visualization, black nodes present the initial subset without information diffusion. White nodes correspond to the examples that are not influenced by diffusion. For other nodes, darker nodes represent earlier influenced examples. We can observe that the subset with high influence (b) can achieve better performance by influencing a larger group of examples in the unlabeled data pool compared to the subset with low influence (a).
@@
To reduce the annotation cost, the previous effort Vote-$k$ (Su et al., 2023) made attempts by proposing to select a diverse and representative subset from a large-scale unlabeled data pool to annotate. Particularly, Vote-$k$ initially selects a small portion of data for diversity and annotates them manually. Then, these annotated data act as prompts for predictions on all other unlabeled data, and choose the remaining ones that need to be annotated, based on diverse confidence scores. However, despite its strong performance in empirical evaluations, Vote-$k$ is still unsatisfactory in practice. We detail the issues from three aspects. (1) The data selection procedure of Vote-$k$ is not end-to-end. This results in inconvenience, increased processing complexity, and added inference costs due to the predictions on unlabeled data. (2) Diversity and representativeness need to be balanced carefully (Su et al., 2023). Highlighting diversity in data selection is crucial for comprehensive coverage, but may sacrifice representativeness by overlooking exemplary data. Besides, the excessive emphasis on diversity of Vote-$k$ causes the selection of outliers (see evidence in Appendix C.2). (3) Vote-$k$ lacks theoretical guarantees, making it challenging to assess the algorithm’s reliability in realistic tasks and constraining its practical utility.
@@
In this paper, to minimize annotation costs for ICL and address the issues of existing work, an innovative data selection method is introduced, where we utilize influence-driven selective annotations to empower in-context learners (IDEAL). In essence, IDEAL aims to identify a subset of data that acts as a proxy and closely approximates the vast unlabeled dataset. Once annotated, these selected data can be considered a viable substitute for the large annotated examples in subsequent ICL tasks. In further detail, our method works in an unsupervised and end-to-end manner. We first construct a directed graph, where its vertices represent unlabeled data and its edges bridge different data based on their similarities. Inspired by influence maximization that aims to select a vertex set at key positions in social graphs (Li et al., 2018), we then propose to quantify the influence of each candidate unlabeled subset in our constructed graph, through a classic independent-cascade diffusion model illustrated in Figure 2. To find the subset with high influence, a simple greedy algorithm for unlabeled data selection is introduced. The algorithm does not need a delicate trade-off between diversity and representativeness. Instead, it iteratively selects a vertex if it provides a maximum marginal gain to the influence metric, until the selection is completed based on the annotation budget.
@@
Theoretically, under the influence-driven selective paradigm, we provide the lower bound for the subset influence selected by our method, demonstrating it is at least as large as a certain proportion of the influence of the optimal solution. Empirically, we conduct comprehensive experiments over 9 datasets across diverse tasks (covering classification, commonsense reasoning, dialogue, and text/code generation). Various LLMs and prompt retrieval technologies are included in evaluations. Experimental results demonstrate that our IDEAL can achieve better performance than Vote-$k$ in 17 out of 18 cases in the experiments, with only 13% time consumption during subset selection. This creates a strong baseline of selective annotations for follow-up research. Source codes have been attached for the reproducibility of results.
@@ 2.2 Influence-driven selective annotations
Overview. For selective annotations in ICL, we need to identify a subset that approximates vast unlabeled data. Therefore, quantifying the coverage of each candidate subset is critical. To achieve this, we construct a directed graph using the embeddings of unlabeled data and portray their relationships using the edges in the graph. We then quantify the influence of each candidate subset in the constructed graph. An information diffusion model is used for this purpose. Through the information diffusion model to quantifying the influence of each candidate subset, we avoid the delicate trade-off between diversity and representativeness. After the quantification, we can search the subset with maximum influence, which most closely approximates the unlabeled data. Below we detail the above procedure step by step.
Constructing the directed graph. We first compute a vector embedding for each unlabeled instance using Sentence-BERT (Reimers & Gurevych, 2019) [^1] https://huggingface.co/sentence-transformers/all-mpnet-base-v2.. The obtained embeddings are employed to build a directed graph $\mathcal{G}=(\mathbf{V},\mathbf{E},\mathbf{P})$, where the vertices $\mathbf{V}=\{\mathbf{v}_{i}\}_{i=1}^{n}$ represent the embeddings of the unlabeled instances, $\mathbf{E}$ denotes the set of edges in the graph, and $\mathbf{P}$ denotes the set of weights assigned to edges. In more detail, for each vertex $\mathbf{v}\in\mathbf{V}$, we connect it to its $k$ nearest successors [^2] In graph theory (Harary, 2018), a vertex $\mathbf{u}$ is the successor of a vertex $\mathbf{v}$ if it is at the end of an outgoing directed edge ($\mathbf{v}$,$\mathbf{u}$). in terms of the cosine similarity between the embeddings and then get $\mathbf{E}$. For the edge $(\mathbf{v},\mathbf{u})\in\mathbf{E}$ that connects $\mathbf{v}$ and its successor $\mathbf{u}$, we assign the weight $p(\mathbf{v},\mathbf{u})=\cos(\mathbf{v},\mathbf{u})/{\sum_{\mathbf{z}\in\mathcal{N}(\mathbf{v},k)}\cos(\mathbf{v},\mathbf{z})}$ with $p\in\mathbf{P}$, where $\mathcal{N}(\mathbf{v},k)$ represents the set including $k$ nearest successors of $\mathbf{v}$, and $\cos(\cdot,\cdot)$ is a function that calculates the cosine similarity of two embeddings. The constructed graph depicts the relationships between unlabeled examples in terms of the embedding similarity.
@@
At the beginning, the subset $\mathcal{S}$ is activated. Then at each step, each vertex $\mathbf{v}$ activates its successors that remained inactive in the last step with a probability defined by $p(\mathbf{v},\mathbf{u})$. The activation can be conceptualized as a coin toss where the outcome is determined by the head probability $p(\mathbf{v},\mathbf{u})$. If the result is the head, the vertex $\mathbf{v}$ becomes activated; otherwise, it remains inactive. Starting from $\mathcal{S}$, the diffusion terminates when no further vertex can be activated in the graph. Lastly, we quantify the influence of the set with the number of activated vertices, where a larger number corresponds to greater influence. In order to get a stable result, we repeat this process ten times and take the average influence. To help understand the procedure of Algorithm 1, we provide an illustration as shown in Figure 2. For convenience, we express Algorithm 1 as an influence function $f_{\mathcal{G}}(\mathcal{S})$ for the graph $\mathcal{G}$ that takes example set $\mathcal{S}$ as inputs, and returns the number of activated vertices $L$.
@@
Searching the subset with maximum influence. We exploit a simple yet effective greedy algorithm (Kempe et al., 2003) to search the subset with maximum influence, which is illustrated in Algorithm 2. Specifically, the algorithm is initialized with an empty set, and iteratively involves an instance if it can provide the maximum marginal gain to the influence function. The search algorithm terminates when the selected subset meets the annotation budget. Finally, we achieve the set $\mathcal{S}_{\rm{u}}$ that includes $m$ examples to annotate, using the correspondence between embeddings and instances. It is worth mentioning that this searching process aims to maximize the influence of the whole selected subset rather than considering each example separately. This is because combining all the individual high-impact examples together does not necessarily achieve the highest-impact subset.
@@ 2.3 Prompt retrieval
After the above influence-driven selective annotations, the subset $\mathcal{S}_{\rm{u}}$ is achieved. By making manual annotations on $\mathcal{S}_{\rm{u}}$, a set of annotated examples is obtained. We can then retrieve examples from the annotated set as in-context examples for each test input. Following previous studies (Liu et al., 2021; Su et al., 2023), we will calculate embeddings for all annotated examples using Sentence-BERT (Reimers & Gurevych, 2019) and identify the most similar instances to each test input based on the cosine similarity. Notice that, the proposed method is agnostic to prompt retrieval methods. As demonstrated in §4.3.3, our method can be combined with any other prompt retrieval technologies. Better prompt retrieval technologies can further boost final performance.
@@ 3 Theoretical Analysis
Remark 1. Intuitively speaking, given any graph $\mathcal{G}$, we say the influence function $f_{\mathcal{G}}$ satisfies the submodular condition if adding one data point to a smaller subset provides more influence than adding the same data point to a larger subset. In other words, it reflects the principle of diminishing returns: the marginal gain of including a data point in a set decreases as the size of the set increases. This condition can hold within the influence function (Li et al., 2019). Considering an extreme case, when subset $\mathcal{S}=\mathbf{V}$, the influence improvement of adding any data point to $\mathcal{S}$ will be zero.
@@ 4.1 Experimental setups
Datasets and tasks. Following previous work (Su et al., 2023), we employ 9 datasets for the evaluations, which can be categorized into 4 different tasks, including classification, multi-choice, dialogue, and generation. The details of the datasets are provided in Appendix D.1. For each dataset, the original “train/dev/test” split from the Transformer library (Wolf et al., 2019) is utilized. We use test data for evaluation if they are available publicly (SST-5 (Socher et al., 2013), DBpedia(Lehmann et al., 2015), MWoZ (Budzianowski et al., 2018), and Xsum (Narayan et al., 2018)). Otherwise, we follow the same setting in (Su et al., 2023) and use the development set. We use accuracy as metric for all classifications and multiple choices tasks, joint accuracy (Budzianowski et al., 2018) for MWoZ, test suite accuracy (Zhong et al., 2020) for GeoQuery (Zelle & Mooney, 1996), and ROUGE-L (Lin, 2004) for Xsum.
Models. If not otherwise specified, we run all experiments on the GPT-J 6B model (Wang & Komatsuzaki, 2021) except the GeoQuery and MWoZ datasets where we use Text-devinci-002 (Chen et al., 2021). We also provide experiments on other models including GPT-Neo 2.7B (Black et al., 2021) and more advanced models GPT-3.5-Turbo (Openai, 2023) in §4.3.4. Our implementation is detailed in Appendix D.2.
Baselines. In the main experiments, we perform a comprehensive evaluation of our method that is compared with previous state-of-the-art selective annotation baselines, i.e., Vote-$k$ (Su et al., 2023) and random selection (abbreviated as “Random” below). Note that, in §4.3.2, we also compare our method with alternative methods that can select a coreset from large-scale unlabeled data on typical datasets. For the baseline Vote-$k$, we conduct experiments by running its official code [^3] https://github.com/HKUNLP/icl-selective-annotation..
@@ 4.2 Main results
Measurement on performance. We first perform the evaluations for Random, Vote-$k$, and our method. The annotation budget is set to 18 and 100 respectively following the same setting as Vote-$k$. Note that we include $18$ as the annotation budget considering all annotated examples can be fit to the prompt of the large language models within context limits. Therefore, the prompt retrieve stage can be ignored and the evaluation results can naturally represent the quality of the selected examples. We provide experimental results in Table 1. As can be seen, our method achieves better performance than baselines in most of the evaluation scenarios (17 out of 18). Interestingly, we find that random selection outperforms Vote-$k$ in 3 out of 18 cases. We conjecture that, under some ideal circumstances, the selected subset by random selection can approximate the distribution of full data. If test data follows the same distribution, good performance can be achieved. Note that we also illustrate selected examples and label distributions in selective annotations in Appendix C.1 and Appendix C.4 respectively.
Measurement on time cost. Previous work Vote-$k$ (Su et al., 2023) encompasses generating prediction for most unlabeled data with a set of selected examples as prompts and performs data selection according to the confidence scores of the prediction. However, this process results in large unnecessary costs at inference time. Meanwhile, LLMs are often used as a service and an extra charge will appear with the usage of the token in both the input and output. In Figure 3, we compare the time cost of subset selection in our method against Vote-$k$ on all tasks with the same hardware. The annotation budget is set to 18. We can observe that our method saves a tremendous amount of cost compared to Vote-$k$. Specifically, under the same hardware conditions, IDEAL achieves a 7.8$\times$ lead on average over Vote-$k$. The speed improvement benefits from the fact that the proposed method does not need to perform example selection by generating predictions on a large number of unlabeled examples and is completely unsupervised.